### v1

# This is the title

   Abstract: We formulate p-adic versions of following three: (1) Grothendieck Inequality, (2) Johnson-Lindenstrauss Flattening Lemma, (3) Bourgain-Tzafriri Restricted Invertibility Theorem.
Keywords: Grothendieck Inequality, Johnson-Lindenstrauss Lemma, Bourgain-Tzafriri Restricted Invertibility, p-adic Hilbert space.
Mathematics Subject Classification (2020): 46S10.
1. p-adic Grothendieck Inequality Problem
In 1953, Grothendieck derived the following inequality which is one of the greatest inequalities of all time.
Theorem 1.1.
[2, 10, 17, 11, 1, 7, 16, 14] (Grothendieck Inequality) There is a universal constant $K_{G}$ satisfying the following: For every Hilbert space $\mathcal{H}$ and for all $m,n\in\mathbb{N}$, if a scalar matrix $[a_{j,k}]_{1\leq j\leq m,1\leq k\leq n}$ satisfies
$\left|\sum_{j=1}^{m}\sum_{k=1}^{n}a_{j,k}s_{j}t_{k}\right|\leq\left(\max_{1\leq j\leq m}|s_{j}|\right)\left(\max_{1\leq k\leq n}|t_{k}|\right),\quad\forall s_{j},t_{k}\in\mathbb{C},$
then
$\left|\sum_{j=1}^{m}\sum_{k=1}^{n}a_{j,k}\langle u_{j},v_{k}\rangle\right|\leq K_{G}\left(\max_{1\leq j\leq m}\|u_{j}\|\right)\left(\max_{1\leq k\leq n}\|v_{k}\|\right),\quad\forall u_{j},v_{k}\in\mathcal{H}.$
We formulate a problem which is p-adic analogue of Theorem 1.1. We first recall the definition of p-adic Hilbert space.
Definition 1.2.
[…]
  1. (i)
    If $x\in\mathcal{X}$ is such that $\langle x,y\rangle=0$ for all $y\in\mathcal{X}$, then $x=0$.
  2. (ii)
    $\langle x,y\rangle=\langle y,x\rangle$ for all $x,y\in\mathcal{X}$.
  3. (iii)
    $\langle x,\alpha y+z\rangle=\alpha\langle x,y\rangle+\langle x,z\rangle$ for all $\alpha\in\mathbb{K}$, for all $x,y,z\in\mathcal{X}$.
  4. (iv)
    $|\langle x,y\rangle|\leq\|x\|\|y\|$ for all $x,y\in\mathcal{X}$.
Following is the standard finite dimensional example.
Example 1.3.
[13] Let $p$ be a prime. For $d\in\mathbb{N}$, let $\mathbb{Q}_{p}^{d}$ be the standard p-adic Hilbert space equipped with the inner product
$\langle(a_{j})_{j=1}^{d},(b_{j})_{j=1}^{d}\rangle\coloneqq\sum_{j=1}^{d}a_{j}b_{j},\quad\forall(a_{j})_{j=1}^{d},(b_{j})_{j=1}^{d}\in\mathbb{Q}_{p}^{d}$
and norm
$\|(x_{j})_{j=1}^{d}\|\coloneqq\max_{1\leq j\leq d}|x_{j}|,\quad\forall(x_{j})_{j=1}^{d}\in\mathbb{Q}_{p}^{d}.$
Based on Theorem 1.1, we formulate following problem.
Problem 1.4.
(p-adic Grothendieck Inequality Problem) Let $\mathbb{K}$ be a non-Archimedean valued field. Whether there is a universal constant $K_{\mathbb{K}}$ satisfying the following property: for every p-adic Hilbert space $\mathcal{X}$ over $\mathbb{K}$, for all $m,n\in\mathbb{N}$, if $[a_{j,k}]_{1\leq j\leq m,1\leq k\leq n}\in\mathbb{M}_{m\times n}(\mathbb{K})$ satisfies
$\left|\sum_{j=1}^{m}\sum_{k=1}^{n}a_{j,k}s_{j}t_{k}\right|\leq\left(\max_{1\leq j\leq m}|s_{j}|\right)\left(\max_{1\leq k\leq n}|t_{k}|\right),\quad\forall s_{j},t_{k}\in\mathbb{K},$
then
$\left|\sum_{j=1}^{m}\sum_{k=1}^{n}a_{j,k}\langle u_{j},v_{k}\rangle\right|\leq K_{\mathbb{K}}\left(\max_{1\leq j\leq m}\|u_{j}\|\right)\left(\max_{1\leq k\leq n}\|v_{k}\|\right),\quad\forall u_{j},v_{k}\in\mathcal{X}.$
Problem 1.5.
(p-adic Grothendieck Inequality Problem) Let $\mathbb{K}$ be a non-Archimedean valued field. Whether there is a universal constant $K_{\mathbb{K}}$ satisfying the following property: for every p-adic Hilbert space $\mathcal{X}$ over $\mathbb{K}$, for all $m,n\in\mathbb{N}$, for every matrix $[a_{j,k}]_{1\leq j\leq m,1\leq k\leq n}\in\mathbb{M}_{m\times n}(\mathbb{K})$, for all $u_{1},\dots,u_{m},v_{1},\dots,v_{n}\in\mathcal{X}\setminus\{0\}$, for all $s_{1},\dots,s_{m},t_{1},\dots,t_{n}\in\mathbb{K}\setminus\{0\}$, we have
$\max_{1\leq j\leq m,1\leq k\leq n}\frac{1}{\|u_{j}\|\|v_{k}\|}\left|\sum_{j=1}^{m}\sum_{k=1}^{n}a_{j,k}\langle u_{j},v_{k}\rangle\right|\leq$ $K_{\mathbb{K}}\max_{1\leq j\leq m,1\leq k\leq n}\frac{1}{|s_{j}||t_{k}|}\left|\sum_{j=1}^{m}\sum_{k=1}^{n}a_{j,k}s_{j}t_{k}\right|.$
2. p-adic Johnson-Lindenstrauss Flattening Problem
Johnson and Lindenstrauss showed that points in high dimensional Euclidean space can be mapped to low dimensional Euclidean space with little distortion to distances between every pair of points, using Lipschitz maps [12].
Theorem 2.1.
[15, 12] (Johnson-Lindenstrauss Flattening Lemma) Let $M,N\in\mathbb{N}$ and $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{M}\in\mathbb{R}^{N}$. For each $10<\varepsilon<$ 1, there exists a Lipschitz map $f:\mathbb{R}^{N}\to\mathbb{R}^{m}$ and a real $r>0$ such that
$r(1-\varepsilon)\|\mathbf{x}_{j}-\mathbf{x}_{k}\|\leq\|f(\mathbf{x}_{j})-f(\mathbf{x}_{k})\|\leq r(1+\varepsilon)\|\mathbf{x}_{j}-\mathbf{x}_{k}\|,\quad\forall 1\leq j,k\leq M,$
where
$m=O\left(\frac{\log M}{\varepsilon^{2}}\right).$
Over the period, some improvements of Theorem 2.1 were obtained. We recall them.
Theorem 2.2.
[9, 12] (Johnson-Lindenstrauss Flattening Lemma: Frankl-Maehara form) Let $0<\varepsilon<\frac{1}{2}$ and $M\in\mathbb{N}$. Define
$m(\varepsilon,M)\coloneqq\left\lceil 9\frac{1}{\varepsilon^{2}-\frac{2\varepsilon^{3}}{3}}\log M\right\rceil+1.$
If $M>m(\varepsilon,M)$, then for any $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{M}\in\mathbb{R}^{M}$, there exists a map $f:\{\mathbf{x}_{j}\}_{j=1}^{M}\to\mathbb{R}^{m}$ such that
$(1-\varepsilon)\|\mathbf{x}_{j}-\mathbf{x}_{k}\|^{2}\leq\|f(\mathbf{x}_{j})-f(\mathbf{x}_{k})\|^{2}\leq(1+\varepsilon)\|\mathbf{x}_{j}-\mathbf{x}_{k}\|^{2},\quad\forall 1\leq j,k\leq M.$
Theorem 2.3.
[4, 12] (Johnson-Lindenstrauss Flattening Lemma: Dasgupta-Gupta form) Let $M,N\in\mathbb{N}$ and $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{M}\in\mathbb{R}^{N}$. Let $10<\varepsilon<$ 1. Choose any natural number $m$ such that
$m>4\frac{1}{\frac{\varepsilon^{2}}{2}-\frac{\varepsilon^{3}}{3}}\log M.$
Then there exists a map $f:\mathbb{R}^{N}\to\mathbb{R}^{m}$ such that
$(1-\varepsilon)\|\mathbf{x}_{j}-\mathbf{x}_{k}\|^{2}\leq\|f(\mathbf{x}_{j})-f(\mathbf{x}_{k})\|^{2}\leq(1+\varepsilon)\|\mathbf{x}_{j}-\mathbf{x}_{k}\|^{2},\quad\forall 1\leq j,k\leq M.$
The map $f$ can be found in randomized polynomial time.
Theorem 2.4.
[12, 8] (Johnson-Lindenstrauss Flattening Lemma: matrix form) There is a universal constant $C>0$ satisfying the following. Let $10<\varepsilon<$ 1, $M,N\in\mathbb{N}$ and $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{M}\in\mathbb{R}^{N}$. For each natural number
$m>\frac{C}{\varepsilon^{2}}\log M,$
there exists a matrix $M\in\mathbb{M}_{m\times N}(\mathbb{R})$ such that
$(1-\varepsilon)\|\mathbf{x}_{j}-\mathbf{x}_{k}\|^{2}\leq\|M(\mathbf{x}_{j}-\mathbf{x}_{k})\|^{2}\leq(1+\varepsilon)\|\mathbf{x}_{j}-\mathbf{x}_{k}\|^{2},\quad\forall 1\leq j,k\leq M.$
Based on Theorem 2.4, we formulate following problem.
Problem 2.5.
(p-adic Johnson-Lindenstrauss Flattening Problem) Let $\mathbb{K}$ be a non-Archimedean valued field. What is the best function $\phi:(0,1)\times\mathbb{N}\to(0,\infty)$ satisfying the following: There is a universal constant $C>0$ (which may depend upon $\mathbb{K}$) satisfying the following. Let $10<\varepsilon<$ 1, $M,N\in\mathbb{N}$ and $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{M}\in\mathbb{K}^{N}$. For each natural number
$m>C\phi(\varepsilon,M),$
there exists a matrix $M\in\mathbb{M}_{m\times N}(\mathbb{K})$ such that
$(1-\varepsilon)\|\mathbf{x}_{j}-\mathbf{x}_{k}\|\leq\|M(\mathbf{x}_{j}-\mathbf{x}_{k})\|\leq(1-\varepsilon)\|\mathbf{x}_{j}-\mathbf{x}_{k}\|,\quad\forall 1\leq j,k\leq M.$
3. p-adic Bourgain-Tzafriri Restricted Invertibility Problem
Let $\mathbb{R}^{d}$ be the standard $d$-dimensional Euclidean space with canonical orthonormal basis $\{e_{j}\}_{j=1}^{d}$. Bourgain and Tzafriri in 1987 showed that any matrix can be inverted restrictively with control over norm [3].
Theorem 3.1.
[3] (Bourgain-Tzafriri Restricted Invertibility Theorem) There are universal constants $A>0$, $c>0$ satisfying the following property. If $d\in\mathbb{N}$, and $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is a linear operator with $\|Te_{j}\|=1$, $\forall 1\leq j\leq d$, then there exists a subset $\sigma\subseteq\{1,\dots,d\}$ of cardinality
$\operatorname{Card}(\sigma)\geq\frac{cd}{\|T\|^{2}}$
such that
$\left\|\sum_{j\in\sigma}a_{j}Te_{j}\right\|^{2}\geq A\sum_{j\in\sigma}|a_{j}|^{2},\quad\forall a_{j}\in\mathbb{R},\forall j\in\sigma.$
Recently, Spielman and Srivastava gave a simple proof of Theorem 3.1 by proving a generalization of Theorem 3.1 [18] due to Vershynin [19]. Based on Theorem 3.1, we formulate following problem.
Problem 3.2.
(p-adic Bourgain-Tzafriri Restricted Invertibility Problem) Let $\mathbb{K}$ be a non-Archimedean valued field. Whether there are universal constants $A>0$, $c>0$ (which may depend upon $\mathbb{K}$) satisfying the following property: If $d\in\mathbb{N}$, and $T:\mathbb{K}^{d}\to\mathbb{K}^{d}$ is a linear operator with $\|Te_{j}\|=1$, $\forall 1\leq j\leq d$, then there exists a subset $\sigma\subseteq\{1,\dots,d\}$ of cardinality
$\operatorname{Card}(\sigma)\geq\frac{cd}{\|T\|^{2}}$
such that
$\left\|\sum_{j\in\sigma}a_{j}Te_{j}\right\|^{2}\geq A\max_{j\in\sigma}|a_{j}|^{2},\quad\forall a_{j}\in\mathbb{K},\forall j\in\sigma.$